OpenStudy (anonymous):

A bag contains 4 red marbles and 6 blue marbles. Brandon selects one marble from the bag. Without replacing the marble he then selects a second marble. What is the probability that he first selects a blue marble and then selects a red marble?

OpenStudy (anonymous):

blue marble first time: 6 blue marbles/10total marbles = 3/5 red marble second time: 4 red marbles/9total marbles = 4/9 Total probability = multiple both probabilities = 4/15